# No Bells, Just Whistles: Sports Field Registration by Leveraging Geometric Properties

### v2

@@ Abstract
Broadcast sports field registration is traditionally addressed as a homography estimation task, mapping the visible image area to a planar field model, predominantly focusing on the main camera shot. Addressing the shortcomings of previous approaches, we propose a novel calibration pipeline enabling camera calibration using a 3D soccer field model and extending the process to assess the multiple-view nature of broadcast videos. Our approach begins with a keypoint generation pipeline derived from SoccerNet dataset annotations, leveraging the geometric properties of the court. Subsequently, we execute classical camera calibration through DLT algorithm in a minimalist fashion, without further refinement. Through extensive experimentation on real-world soccer broadcast datasets such as SoccerNet-Calibration, WorldCup 2014 and TS-WorldCup, our method demonstrates superior performance in both multiple- and single-view 3D camera calibration while maintaining competitive results in homography estimation compared to state-of-the-art techniques.  [^1] https://github.com/mguti97/No-Bells-Just-Whistles
1 Introduction
@@
Figure 1: Overview of our proposed framework. Top: Training data generation pipeline. Beginning with SoccerNet [8] annotations, we utilize field line extraction and ellipse fitting to establish a hierarchical structure for computing each set of keypoints. Bottom: The encoder-decoder networks produce heatmaps for keypoints and extremities of soccer field lines to extract their positions in the image space. The obtained keypoint set is augmented with intersections of lines generated by the second model to ensure a sufficient number of points.
Camera calibration is essential for a wide range of computer vision applications, such as structure from motion [13, 14, 4, 2], reconstruction [26, 3], and pose estimation [9, 38]. In the sports domain, accurately estimating pairwise correspondences between the sports field and broadcast video is crucial for sorting out some high-level tasks. This process not only streamlines manual labor but also enhances the visual appeal of broadcast matches through augmented reality and virtual advertisement insertion, while also facilitating the development of advanced tools for sports analytics.
@@
Sports fields, with their well-defined dimensions [1], serve as calibration objects. However, achieving accurate camera calibration in the broadcast setting poses challenges due to multiple camera views and partial occlusion of the court, hindering the matching process between 2D and 3D correspondences. Additionally, the variability of the camera focal length further complicates calibration efforts. The recent surge in deep learning has led to several data-driven approaches for field-specific feature prediction [9, 27, 7, 28, 21] or direct homography matrix regression [23, 32]. Others investigate camera calibration as a search problem [30, 6, 29, 36, 37], generating camera pose databases and refining homography estimates to improve calibration accuracy. Moreover, some approaches [18, 17, 9, 27, 10] leverage temporal homography consistency between video frames, intending to better align with the nature of sports video broadcasts. Focusing on the soccer domain, despite the potential of estimating camera parameters for reconstructing non-planar points and enabling applications such as automatic camera control, offside detection or 3D ball tracking, previous studies [6, 29, 30, 23, 9, 36, 37, 27, 31, 7, 28] have predominantly treated the task as homography estimation rather than full calibration [33]. Inspired by the limitations of existing approaches, we propose a novel geometry-based keypoint retrieval pipeline (see Figs. 1-2) for 3D sports field registration, additionally capable of addressing the challenges posed by the multiple-view broadcast nature. This approach involves defining a hierarchical pipeline to extract a pre-defined keypoint grid from the court’s geometric properties and leveraging an encoder-decoder neuronal to estimate keypoint positions. Specifically, we adopt HRNetv2 [34] as the backbone model for the keypoints prediction. Finally, the estimated keypoints are used to compute the projection matrix using RANSAC [16] and Direct Linear Transformation (DLT) [20] algorithms. The pipeline of our proposed method is outlined in Fig. 1. We extensively evaluate our approach on three real-world soccer broadcast datasets, (SoccerNet-Calibration [8], WorldCup 2014 [21] and TS-WorldCup [7]) and compare it with state-of-the-art methods in both 2D and 3D sports field registration. The experiments demonstrate that our model achieves superior performance on 3D camera calibration while maintaining comparable results on homography estimation with respect to competing approaches.
In a nutshell, this paper makes the following contributions:
@@
  • A pipeline structure focused solely on 2D-3D correspondences, without further refinement. In other words, a minimalist approach without bells and whistles—or, in soccer terms, no bells, just whistles.
2 Related Work
@@
Sports field registration is a critical component of most sports applications in computer vision, whose common approaches intend to estimate homography matrices in team sports. Homography estimation has traditionally relied on the corresponding features, or keypoints, identification between images and the court field model. These features, usually obtained by exploiting geometric primitives such as lines and/or circles, are subsequently used to estimate the mapping between the images by using RANSAC [16] algorithm with DLT [20] or non-linear optimization through the minimisation of a chosen loss function. More recent approaches either directly predict an initial homography matrix. Alternatively, they seek the optimal matching homography within a reference database that includes synthetic images with known homography matrices.
Homography Estimation. Various recent methods for sports field registration obtain an initial homography estimation through a grid of uniformly sampled and predicted keypoints [9, 27, 7, 24, 22, 10, 28], or line and circle pixels using semantic segmentation [21, 29, 32, 36, 37, 33]. These are obtained with Deep Neural Networks (DNN) and used as corresponding features between the field template and camera image. Following a prediction-based strategy, [23, 33] proposed an optimization-based framework to obtain homography estimation and camera parameters by minimizing image registration and segment reprojection error, respectively. Shi et al. [31] proposed an iterative estimation process to estimate any homography transformation, regardless of the degree of misalignment between the image and the template, in a self-supervised manner. Moreover, end-to-end approaches [32] have been proposed with promising results. Search-based approaches involve creating databases and searching for the best matching homography through the usage of edge maps [30, 6] or semantic segmentation [29, 36, 37]. However, while search-based methods are highly accurate and robust, they often incur a significant computational cost due to their time-consuming processing steps. Furthermore, recent approaches [9, 27, 10] leverage broadcast sequentiality to enforce temporal consistency between subsequent frames’ homographies.
Homography Refinement. Homography refinement is a crucial step in camera calibration, aiming to achieve an even more accurate homography estimation, if necessary. Previous works use one or a combination of the following methods to refine the initial estimate. [23] and [31] directly regress the refined homography from the input image and field model with DNNs. [6, 30] obtained an estimate for the camera pose by matching with a feature-pose database, and used 100k templates to ensure the assumption of small transform between the input image and the matched template. Iterative optimisation of the camera pose or homography is also proposed in previous works [33, 6, 23], commonly based on re-projection or registration error. Moreover, other approaches [29, 27] exploit the temporal consistency between subsequent video frames.
Camera Parameters Retrieval. Approaches using feature-pose synthetic databases [30, 6, 29, 36, 37] allows for homography and/or direct camera parameters retrieval as templates are created through projective geometry. However, due to small transform assumption, smaller databases lead to larger reprojection errors, making homography refinement a crucial step. Carr et al. [5] estimate camera’s extrinsic and intrinsic parameters leveraging gradient-based alignment to edge images. Moreover, homography decomposition [20] allows for access to individual camera parameters, as shown in [9, 33].
3 Methodology
@@
Sports TV broadcasts consist of sequences of images featuring a fraction of the sports field from an uncalibrated moving camera perspective. Our approach focuses on both extrinsic and intrinsic camera parameters retrieval from each individual frame without any prior information about the camera position or orientation. The proposed method comprises four processing components: Soccer field modelling and keypoint generation, keypoint and line detection, DLT algorithm, and camera parameters retrieval. Next, these components are introduced.
3.1 Modelling the Soccer Field
@@
A soccer field is composed of lines and circle segments, representing all field markings, goal posts, and crossbars. Our approach, like keypoint-based methods [32, 9, 27, 7, 24, 10, 28, 15], relies on the lines painted on the ground, their intersections, and the corners they define, due to its known position on the world coordinate system. We follow the segment definitions of Cioppa et al. [8], and set them as starting points to hierarchically compute our pre-defined keypoints from court geometric properties.
3.1.1 Keypoint Generation
@@
The full set of sampled keypoints is organized into subgroups based on the specific geometric features they represent (Fig. 2). The hierarchical nature of the computation ensures that information from initially identified keypoints is exploited for computing the subsequent ones (some instances in Fig. 1-top). Next, we define the keypoints sets:
@@
  • Line-Line intersections. Following [32, 11, 9, 10, 28, 15], this set of points ($\mathcal{K}p$) includes the intersections of boundary lines or the penalty area markings. Considering the 23 lines depicted in [8], including goal posts and crossbars, up to 30 points can be included.
@@
  • Extended Line-Line intersections. Following [11], this set ($\mathcal{K}e$) addresses the intersections of extended lines that represent non-adjacent segments of the soccer field. To obtain that, the lines are extended by exploiting their equations beyond their original boundaries.
@@
  • Line-Ellipse intersections. Following [32, 11, 9, 10, 28, 15], this set ($\mathcal{K}p_{1}$) is to consider the intersections between the field lines and the circles or semi-circles present on the court. Given the distortions introduced by the camera perspective, conics on the field are considered ellipses for equation computation. The parameters of these ellipses are fitted using the least squares method [19]. Intersection points were analytically derived using ellipse and line formulas.
@@
  • Ellipse tangent points. Following [18, 15], augmentation of available points is achieved through the utilization of tangent points on tangent lines, extending from a specified point to the previously defined ellipses. These tangent points (denoted by $\mathcal{K}p_{2}$) were analytically determined by employing an ellipse equation, incorporating the known coordinates of an external point.
@@
  • Additional points. Following [7, 10, 24, 27, 28, 15], once the previous points and the corresponding homography are inferred, an additional set ($\mathcal{K}p_{3}$) of 9 points is integrated along the central pitch axis, encompassing the pitch center and penalty points. Additionally, 4 points are strategically placed to designate quarter turns along the central circle. Furthermore, the homography facilitates the inclusion of other points that are initially missing, addressing situations such as unannotated lines.
3.1.2 Keypoint Disambiguation

### v3

# PnLCalib: Sports Field Registration via Points and Lines Optimization

@@ Abstract
Camera calibration in broadcast sports videos presents numerous challenges for accurate sports field registration due to multiple camera angles, varying camera parameters, and frequent occlusions of the field. Traditional search-based methods depend on initial camera pose estimates, which can struggle in non-standard positions and dynamic environments. In response, we propose an optimization-based calibration pipeline that leverages a 3D soccer field model and a predefined set of keypoints to overcome these limitations. Our method also introduces a novel refinement module that improves initial calibration by using detected field lines in a non-linear optimization process. This approach outperforms existing techniques in both multi-view and single-view 3D camera calibration tasks, while maintaining competitive performance in homography estimation. Extensive experimentation on real-world soccer datasets, including SoccerNet-Calibration, WorldCup 2014, and TS-WorldCup, highlights the robustness and accuracy of our method across diverse broadcast scenarios. Our approach offers significant improvements in camera calibration precision and reliability.  [^1] https://https://github.com/mguti97/PnLCalib
Index Terms:
Camera Calibration, Homography Estimation, Sports Analytics, SoccerNet, World Cup.
I Introduction
Sports analytics has become an increasingly vital component in modern sports, transforming the way teams, coaches, and fans understand and optimize athletic performance. The proliferation of advanced tracking technologies, such as player and ball tracking systems, has enabled the generation of rich, high-resolution data that provides unprecedented insights into the dynamics of sports competitions. This wealth of tracking data has revolutionized the way sports are analyzed, allowing for more informed decision-making, enhanced player development, and the identification of strategic advantages. Leveraging these data-driven insights has become a key competitive edge, as teams strive to gain a deeper understanding of player movements, team tactics, and in-game patterns. The ability to accurately capture and analyze tracking data has become a crucial aspect of sports analytics, fueling innovations in areas like player performance optimization, injury prevention [4], and the development of advanced coaching strategies [51].
While the proliferation of wearable tracking devices has been instrumental in generating sports performance data, the use of computer vision techniques has emerged as a compelling alternative approach. By leveraging advanced computer vision algorithms, researchers and sports organizations can now extract valuable tracking data directly from video footage, without the need for intrusive wearable sensors [35]. This camera-based tracking approach offers several advantages, including the ability to capture data from multiple athletes simultaneously, the elimination of potential interference or disconnection issues associated with wearables, and the potential for retroactive analysis of historical game footage. Computer vision-based tracking leverages techniques such as object detection, object tracking, and pose estimation to accurately identify and monitor the movements of players, balls, and other key elements within a sports environment. This data-driven, non-invasive approach to tracking has become increasingly sophisticated, enabling the generation of rich, high-fidelity datasets that can provide deeper insights into athletic performance and team dynamics.
The applications of computer vision in sports extend far beyond just tracking player and ball motions. Innovative computer vision tools have been leveraged to enhance various aspects of the sports experience. One prominent example is the use of semi-automatic offside detection systems, which leverage computer vision algorithms to quickly and accurately determine offside positions during live matches, providing crucial support to referees and improving the fairness and pace of the game. Additionally, computer vision techniques are being employed to generate real-time graphics and overlays for sports broadcasts, seamlessly integrating information such as player statistics, team formations, and tactical visualizations. This enhanced visual experience not only informs and engages the audience but also creates new opportunities for data-driven storytelling and fan engagement. Looking ahead, the continued advancements in computer vision are poised to revolutionize various facets of sports, from automated refereeing and in-depth performance analytics to immersive fan experiences and the integration of augmented reality into the viewing experience.
While the advancements in sports analytics and computer vision have enabled unprecedented insights and experiences, one crucial aspect that underpins these capabilities is the accurate calibration of cameras used to capture sports footage. Camera calibration refers to the process of determining both intrinsic and extrinsic parameters of a camera system, which is essential for transforming 2D image data into meaningful 3D representations of the sports environment. While sports fields, with their well-defined dimensions [42], serve as calibration objects, achieving accurate camera calibration in the broadcast setting poses challenges due to multiple camera views, focal length variability and partial occlusion of the court, hindering the matching process between 2D and 3D correspondences.
@@ I Introduction
Figure 1: Overview of our proposed framework. Top: Training data generation pipeline. Beginning with SoccerNet [8] annotations, we utilize field line extraction and ellipse fitting to establish a hierarchical structure for computing each set of keypoints. Bottom: Inference stage pipeline. The encoder-decoder networks produce heatmaps for keypoints and extremities of soccer field lines to extract their positions in the image space. The obtained keypoint set is augmented with intersections of lines generated by the second model to ensure a sufficient number of points. After initial calibration, our PnL refinement module is applied to further refine the calibration estimate by jointly using detected points and lines information.
@@
Traditional sports field registration relied on feature-based methods [37], such as detecting and matching local features like SIFT [27] and MSER [31] to estimate pairwise correspondences and compute a homography matrix by using RANSAC [15]. The recent surge in deep learning has led to several data-driven approaches leveraging Convolutional Neural Networks (CNNs) for feature extraction, showing promising results in sports field registration. These methods include field-specific feature prediction [9, 33, 7, 34, 23, 14, 19] and direct homography matrix regression [26, 45]. Other researchers have investigated camera calibration as a search problem [39, 5, 38, 52, 53, 32], generating camera pose databases and refining estimates to improve calibration accuracy. Moreover, some approaches [2, 18, 16, 9, 33, 10] leverage temporal calibration consistency between video frames, intending to better align with the nature of sports video broadcasts. Focusing on the soccer domain, despite the potential of estimating camera parameters for reconstructing non-planar points and enabling applications such as automatic camera control, offside detection, or 3D ball tracking, previous studies [5, 38, 39, 26, 9, 52, 53, 33, 40, 7, 34] have predominantly treated the task as homography estimation rather than full calibration [46, 32].
Inspired by the limitations of existing approaches, we propose a novel calibration pipeline (see Fig. 1) for 3D sports field registration. An early version of this work was presented in [19], in which we proposed our method to be capable of addressing the challenges posed by the multiple-view broadcast nature. This approach involves defining a hierarchical pipeline to extract a pre-defined keypoint grid from the court’s geometric properties and leveraging an encoder-decoder network to estimate keypoint positions. Moreover, the soccer field’s lines are defined following the SoccerNet [8] notation and line extremities are also extracted. Particularly, we adopt HRNetv2 [49] as the backbone model for the keypoints and line extremities prediction. The estimated keypoints are used to compute an initial estimate of the projection matrix using RANSAC [15] and Direct Linear Transformation (DLT) [22] algorithms. In this paper, we extend our contribution by incorporating a novel refinement module that jointly uses the detected keypoints and lines further to optimize the initial estimate as a non-linear least-squares problem [47]. We extensively evaluate our approach on three real-world soccer broadcast datasets, including SoccerNet-Calibration [8], WorldCup 2014 [23], and TS-WorldCup [7] datasets, and compare it with state-of-the-art methods in both 2D and 3D sports field registration. The experiments demonstrate that our model achieves superior performance on 3D camera calibration while maintaining comparable results on homography estimation with respect to competing approaches. In summary, this paper makes the following contributions:
@@
  • A refinement module able to optimize the calibration estimate by jointly using the detected keypoints and lines.
II Related work
@@
Sports field registration is a critical component of most sports applications in computer vision, whose common approaches intend to estimate homography matrices in team sports. Traditionally, homography estimation has relied on identifying corresponding features or keypoints between images and the court field model. These features, typically obtained by exploiting geometric primitives such as lines and/or circles, are subsequently used to estimate the mapping between the images. This is often done using the RANSAC algorithm [15] in conjunction with DLT [22] or non-linear optimization techniques [47] that minimize a particular loss function. More recent approaches have diverged from this traditional method. Some directly predict an initial homography matrix, while others seek the optimal matching homography within a reference database containing synthetic images with known homography matrices. Furthermore, the latest approaches have shifted towards directly retrieving camera parameters instead of the homography matrix. This is achieved through various means, including direct prediction, optimization techniques, leveraging databases of image-pose pairs, or decomposing the homography matrix.
II-A Search-based Methods
A prevalent approach in the field has been generating synthetic data to populate databases with homographies or camera poses paired with corresponding image features. These features are often derived from edge maps or semantically segmented images, which represent key elements of the sports field such as lines, circles, and other distinctive markings. Sharma et al. [39] created a synthetic database of edge map-homography pairs. One of the main drawbacks is using normal distribution in the camera pose sampling, which often leads to non-realistic poses. Addressing this issue, Chen and Little [5] created a features-pose database by deducing statistics derived from WorldCup 2014 dataset [23]. Those statistics, related to the camera parameters, were used to sample 90,000 poses which will be encoded through a Siamese Network [20] to distinguish different edge maps. During inference, the network extracts the encoding from edge maps and searches database for nearest neighbour candidate pose. By using area-based semantic segmentation of the soccer field instead of edge images, in contrast with previous approaches, Sha et al. [39] also generated an artificial camera pose database based on the possible ranges for pan and tilt angles and focal length parameters. Aiming to solve image-to-image translation problems, Zhang et al. [52] use improved semantic segmentation using a conditional generative adversarial network [24]. Overall, search-based methods for camera calibration face inherent trade-offs between database size, processing speed, and estimation accuracy. Smaller databases offer faster searches but may compromise initial estimation quality, while larger databases provide better accuracy at the cost of increased computational time. Additionally, these methods often struggle with non-common camera poses frequently encountered in broadcast videos, such as close-ups or oblique angles, which are typically underrepresented in predefined databases.
II-B Optimization-based Methods
Alternatively, common approaches also use edge maps, semantically segmented images or information extracted from the field’s visual landmarks, like intersection points or lines, to obtain homography matrix or camera parameters from optimization methods. Homayounfar et al. [23] classifies field lines using a modified VGG network [41] and extracts their vanishing points reducing the effective number of degrees of freedom (DoF) of the homography from 8 to 4. The field localization problem is formulated as a Markov random field and it requires at least a pair of both vertical and horizontal lines to estimate the vanishing points. Citraro et al. [9] make use of a U-Net architecture [9] to jointly detect semantic keypoints corresponding to field’s line intersections and player positions. In addition to obtaining the homography matrix from the 2D-3D correspondences, intrinsic and extrinsic camera parameters are subsequently obtained through homography decomposition [22]. In order to alleviate the problem of visual landmarks sparsity, Nie et al. [33] propose an encoder-decoder network that jointly outputs a grid of keypoints distributed uniformly on the field template and dense template-features in order to further refine the initial keypoint based homography estimate. Similarly, Maglo et al. [30] use an encoder-decoder network to predict a perspective-aware keypoint grid. This is, the points nearest to the camera are more spread out than the point farthest to the camera in order to compensate for the too large distance variations in the image generated by the perspective effect. More recently, Chu et al. [7] also use a grid of 91 uniformly distributed keypoints and formulate its detection problem as an instance segmentation with dynamic filter learning. This is, the convolution filters are generated dynamically, conditioned on the field image and associated keypoint identity. Oo et al. [34] use a Residual EfficientNet-Attention UNet architecture to estimate the initial homography matrix using pre-defined keypoints to register sports fields. The encoder uses the EfficientNetV2 [44] as the backbone network, and the decoder consists of deconvolution layers with residual blocks, skip connections and attention gates. Theiner et al. [46] introduce a differentiable objective function that is able to learn the camera pose and focal length from segment correspondences. Instance segmentation for each visible line or circle segment is achieved with ResNet [6] backbone and then, the segment reprojection error induced by the estimated camera parameters is iteratively minimized with a gradient-based method [1]. Finally, leveraging the temporal consistency from sports videos, Claasen et al. [10] propose a Bayesian framework. Inspired by recent developments in tracking-by-detection methods, this work proposes a dynamics model that explicitly relates image keypoint positions from one frame to the next through two stages: the first stage consists of a linear Kalman filter, which considers the image keypoints the only part of its state vector, and the second stage incorporates the initial homography estimate to an Extended Kalman Filter [25], with the assumption that the relative homography between frames is small. Falaleev et al. [14] significantly increases the number of usable points for calibration by exploiting line-line and line-conic intersections, points on the conics, and other geometric features, followed by a DLT optimization. Although the computational cost of the enumerated approaches is lower, optimizion-based methods are, overall, not as robust and accurate as search-based methods. Moreover, optimization-based methods are heavily dependent on the landmark detection accuracy.
II-C Prediction-based Methods
Prediction-based approaches use Deep Neural Network-based (DNN) models to calibrate the moving camera in sports. Jiang et al. [26] train a DNN that directly regresses a homography parameterization given an input frame. Subsequently, a sports field template is warped according to the initial estimate. A concatenation of the input frame and the warped template is fed into a second DNN that estimates the error of the current warping and the estimated homography is accordingly optimized until convergence. Tarashima et al. [45] propose SFLNet, a CNN single shot regressor that jointly predicts several outputs: a court metric model defined as an 8-dimensional parameter set, which correspond to the homography’s DoF, a court semantic segmentation which divides the sports field’s spatial layout into divide a frame into court, person, and background regions, and a label adjacency, which comprises adjacencies of label pairs in addition to their presence in an input frame, regularizing the model training via exploiting contextual information. Towards the design of an end-to-end approach, Shi et al. [40] propose a self-supervised learning method for homography estimation. This work employs a self-supervised data mining method to train the registration network with an image and its edge map by using an iterative estimation process controlled by a score regression network to measure the registration error. This method is able to obtain competitive results with previous approaches without the need of any labelled data. Recently, Zhang et al. [53] proposed a four-point calibration method. A cGAN is used to generate semantically segmented frames, eliminating foreground objects. Subsequently, a regression network estimates four points from the frames, which will be used to calculate a homography using DLT algorithm, keeping computational cost low. Lastly, extending to the multiple-view nature of the sports broadcast videos, Mavrogiannis et al. [32] propose a camera calibration method based on synthetic poses. They build a training data generation pipeline with separate flows for the main-camera frames and the rest of camera locations. For the first case, an EfficientNet [6] model is trained to take synthetic edge images as input and regress the camera location. For the second case, a perspective transformation is calculated from four annotated pairs of corresponding points, and the camera location is obtained through homography decomposition. Moreover, a second model is trained on edge maps produced from the poses for each of the possible camera locations to estimate rotation angles and focal length of the camera. Although yielding competitive results on camera calibration and extending it to multiple views, the location-dependent auxiliary model limits the generalization capability and versatility of the method.
II-D Homography and Calibration Refinement
Homography refinement is a crucial step in camera calibration, aiming to achieve an even more accurate homography estimation and camera calibration, when necessary. Previous works use one or a combination of the following methods to refine the initial estimate: Puwein et al. [37] refine the initial homography by bundle adjustment [47]. Chen et al. [5] use the Lucas-Kanade algorithm [3] to refine the initial homography matrix reducing the distance from every pixel in the testing image to its closest edge pixels from a retrieved edge image. Sha et al. [38] introduce the spatial transformer network (STN) to handle large non-affine transformation. The method stacks the input semantic image and the selected template to feed the STN, which outputs a relative homography that maps the semantic image to the template. Citraro et al. [9] use the detected players’ position on the court’s plane to refine the initial homography estimate. After homography decomposition is performed to extract intrinsic and extrinsic camera parameters, these are further refined with Levenberg–Marquardt [22] algorithm. A common approach [52, 53, 32] is to use enhanced correlation coefficient method [13], which performs image alignment between its inputs and returns the refined homography matrix. Alternatively, Oo et al. [34] train an homography refinement network by applying random perturbations to input images. The network outputs the transformation matrix to be applied to the initial estimate. Lastly, other approaches exploit the temporal consistency between subsequent video frames information, such as dense feature maps [33], homography matrices [33, 10] or intrinsic and/or extrinsic camera parameters [9].
III Methodology
@@
Sports TV broadcasts consist of video sequences featuring a fraction of the sports field from different uncalibrated moving camera perspectives, defining a multiple-view setting. Our approach focuses on retrieving both extrinsic and intrinsic camera parameters from each individual frame, without any prior information about the camera’s position or orientation, except for a partial view of the soccer field. The proposed method comprises five processing components: soccer field modelling and keypoints generation, keypoints and line extremities detection, DLT algorithm and camera parameters retrieval, as well as calibration estimate refinement. Next, these components are introduced.
III-A Modelling the Soccer Field
@@
A soccer field is composed of lines, circles and semi-circles, representing all field markings, goal posts, and crossbars. Our approach, following keypoint-based methods [45, 9, 33, 7, 30, 10, 34, 14], relies on the lines painted on the ground, their intersections, the corners they define and some extra geometric properties, due to its known position on the world coordinate system. We follow the segment definitions of Cioppa et al. [8] and set them as starting points to hierarchically compute our pre-defined keypoint grid using court geometric properties.
III-A1 Keypoint Generation
@@
The full set of sampled keypoints is organized into subgroups based on the specific geometric features they represent (see Fig. 2). The hierarchical nature of the keypoint generation pipeline ensures that information from initially identified keypoints is exploited for computing the subsequent ones (some instances in Fig. 1-Top). Next, we define the keypoints sets:
@@
  • Line-Line intersections. Following [45, 11, 9, 10, 34, 14], this set of points ($\mathcal{K}p$) includes the intersections of boundary lines, and the penalty or goal area markings. Considering the 23 lines depicted in [8], including goal posts and crossbars, up to 30 points can be included.
@@
  • Extended Line-Line intersections. Following [11], this set ($\mathcal{K}p_{e}$) addresses the intersections of extended lines that represent non-adjacent segments of the soccer field. To obtain this set, the field lines are extended by exploiting their line equations beyond their original boundaries. It is worth noting that not all non-adjacent line intersections are added; if the lines have to be largely extended, small errors in the line equation would lead to huge deviations in the intersection position.
@@
  • Line-Ellipse intersections. Following [45, 11, 9, 10, 34, 14], this set ($\mathcal{K}p_{1}$) considers the intersections between the field lines and the circles or semi-circles present on the court. Given the distortions introduced by the camera perspective, conics on the field are considered ellipses for its equation computation. The parameters of these ellipses are fitted using the least squares method [21]. Line-ellipse intersection points are analytically derived using ellipse and line formulas (see Fig. 2-Bottom for a visual example).
@@
  • Ellipse tangent points. Following [18, 14], the augmentation of available points is achieved through the utilization of tangent points on tangent lines, extending from a specified external point to the previously defined ellipses. These tangent points (denoted by $\mathcal{K}p_{2}$) were analytically determined by employing an ellipse equation and incorporating the known coordinates of an external point (see Fig. 2-Bottom for a visual example).
@@
  • Additional points. Following [7, 10, 30, 33, 34, 14], once the previous keypoint sets and the corresponding homography are inferred, for the sake of grid completeness, an additional set ($\mathcal{K}p_{3}$) of nine points is integrated along the central pitch axis, encompassing the pitch center and penalty points. Additionally, four points are strategically placed to designate quarter turns along the central circle. Furthermore, the homography facilitates the inclusion of other points that are initially missing, addressing situations such as unannotated or wrongly annotated lines.
III-A2 Keypoint Disambiguation